OpenStudy (anonymous):

p = $14,300 r = 7 1/2% t = 4 What is i?

OpenStudy (lucaz):

find the interest?

OpenStudy (anonymous):

1.) Let x be the number. Two-thirds of this number is (2/3)x and 9 subtracted from it is (2/3)x - 9. This is equal to 7. Solve for x. (2/3)x - 9 = 7 (add 7 to both sides) (2/3)x = 16 (multiply both sides by 3/2) x = 24 <=== 2.) The formula for the circumference of a circle is 2pi(r), where r is the radius. Solve for r. C = 2pi(r) 132 = 2(22/7)r 132 = (44/7)r 21 = r The radius is 21 yards. 3.) Let p be the cost of a pound of potatoes, a be the cost of a pound of apples, and o be the cost of a pound of onions (in cents). Make a system of equations: 9p = 6a (nine pounds of potatoes is the same as six pounds of apples) p = 2o (one pound of potatoes is twice as much as a pound of onions) a = o + 24 (a pound of apples costs 24 cents more than a pound of onions) Solve the system by substitution. 9p = 6a (substitute 2o for p and o + 24 for a) 9(2o) = 6(o + 24) 18o = 6o + 144 12o = 144 o = 12 a = o + 24 = 12 + 24 = 36 p = 2o = 2(12) = 24 A pound of apples costs 36 cents, a pound of onions costs 12 cents, and a pound of potatoes costs 24 cents. 4.) Use the simple interest formula I = Prt = 14300(0.075)(4) = $4290.

OpenStudy (anonymous):

formula is I=(p)(r)(t)

OpenStudy (lucaz):

simple or compound interest?

OpenStudy (anonymous):

Oh so all I do is multiply p r and t? @Philippians413

OpenStudy (anonymous):

s

OpenStudy (anonymous):

yea if it is simple interest that is a simple interest formula that answer should come out as 4,290

OpenStudy (anonymous):

Okay thanks but what is the formula to get p? its on a different question

OpenStudy (anonymous):

p=i/(r)(t) is the formula for finding the principle

OpenStudy (anonymous):

Okay thanks so much for the help :)

OpenStudy (anonymous):

no problem